Given the equation and solution, find the missing coefficient.
8x − 6 = ? x − 7
Solution: x=6

Answer by wilft1(217)

lets call the missing coeffient "y"
8x - 6 = yx-7
substitute what we know for x
8(6) - 6 = y(6) - 7
48 - 6 = y(6) - 7
42 = y(6) - 7
add 7 to both sides
49 = y(6)
divide both sides by 6
y = 8.16